An object of density 300kgm-3 and volume of 0.001m3 is immersed in a water of density 200kgm-3, take g to be equal to 9.8m/sec2.determine;
(a):the weight in air
(b):the weight inside the water
(c):the upward force exerted by the water
(d):relative density
1
Expert's answer
2020-02-25T10:19:17-0500

Given:

"\rho=300\:\rm kg/m^3;"

"V=0.001\:\rm m^3;"

"\rho_W=200\:\rm kg/m^3."


(a) The weight in air

"W_1=mg=g\rho V=9.8\times 300\times 0.001=2.94\:\rm N."

(b) The weight inside the water

"W_2=mg-g\rho_WV=9.8\times 100\times 0.001=0.98\:\rm N."

(c) The upward force exerted by the water

"F_B=g\rho_W V=9.8\times 200\times 0.001=1.96\:\rm N."

(d) The relative density

"\rho/\rho_W=300/200=3/2=1.5"